OpenStudy (anonymous):

To decorate the area around the pool with flowers, your neighbor purchases conical urns of radius 1.5 ft and height 3 ft. How much soil must he buy to fill each urn? a. 2.07 cu. ft per urn c. 8.07 cu. ft per urn b. 7.07 cu. ft per urn d. 27.07 cu. ft per urn

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Do you have ANY diagram?

OpenStudy (anonymous):

not for this one no

OpenStudy (akashdeepdeb):

Haha..You got a pic! :D

OpenStudy (anonymous):

yep :3

OpenStudy (akashdeepdeb):

How many conical uns are there? :D

OpenStudy (anonymous):

idk it doesn't say

OpenStudy (akashdeepdeb):

Okay, I was complicating this question! Just find the Volume of a conical urn = pi * radius * radius * height. = pi * 1.5 * 1.5 * 3 ft^3 Now find the volume of one urn! :D

OpenStudy (anonymous):

i got 190.755 :3

OpenStudy (akashdeepdeb):

3.14 * 2.25 * 3 That CANNOT be 190.755 :P

OpenStudy (anonymous):

21.195 :p

OpenStudy (akashdeepdeb):

I am sorry for that! Colume of a cone = 1/3 * whatever your answer is. = 1/3 * pi * r * r * h Now solve it. You'll get the answer. :)

OpenStudy (anonymous):

thanks got 6.3585 so im thinking B is the closest right?

OpenStudy (akashdeepdeb):

Yes. :)